LION-AG Experiment

  • This is my attempt at transcribing the notes





    LION Test Run

    ALAN Made 3 Test Blocks at L.F.H ONE for Himself. ONE for MFMP-Alan in America has one. ONE was sent to LION.

    My recollection is that they were all calibrated to 1°C accuracy. Alan in America would have done his own thorough calibration run, so both his and Alan's can be relied upon to be believable to sceptics and therefore as a guide to the general characteristics of THIS block.

    When I received the Block from (ALAN LFH) I put a X on the top so to identify it easily from my own several LFH Blocks.

    The Block has not been ran till the day of this test so the glass tubes & coils are in the same condition as when not run by ALAN. The following short calibration is not to test or corroborate ALAN's own test results since I unconditionally accept what he has recorded-stored as data and reported.

    The sole purpose of this calibration is to establish for myself the various PID temperatures, attained when the DIAL of the rheostat/pulser is set to a specific N° on the DIAL.

    When I received it, there was N° dial just the Block knob with its blue line indicator so I made my own. The DIAL is held in Place by locking nut and so is immobilized, it has [????] been superglued. It is this certainty that it can not move that make it possible to accurately and with certainty set a position 1, 63-5 or whatever and therefore know the max current & temp at any setting precisely.

    You will see why clarity on this point is so important, later during the Main Test.

    • The PID thermocouple is connected to the RH Pot containing the Diamond Pads + D2 wrapped in Stainless Steel.
      • T1 = [oo]-TI oo TI = Front Drilled port (Top o)
      • T2 = is from LH Port at the Recv

    [Photo]


    All of the equipment used was from LFH.

    The PSU is a S-400-36 Input is 240-V-AC_ at 4 Amps 36V to 11 AMPS D.C. Output.

    • The PSU with N° LFH Block connected:
      • Draws 10.1 Watts AC falling to 8.5 watts at switch on. This is probably down to the PSU cooling Fan.
    • For my own Measurement purposes I bought and installed a ACM20
      • see (attached) from [RS]-Components. Which was between switches & PSU.
    • DATA Card was provided by ALAN-MARTIN. This New card was used for this test.

    Calibration Run.

    Sunday November 4th 2018

    Start Time - Approx 6.25 pm.

    1. DIAL set to 1
      • Reactor Run empty is N° Tubes in Place
      • Just Thermocouples
      • ACM20 = 13.7 Watts.
      • Ambient 24.5 in EX Room.
      • PID Reachers = 33°C
    2. At 7:30 PM
      • DIAL set to 2.
      • ACM20 = 8.28 Watts
      • PID Reaches = 111°C
    3. At 8:30 PM
      • DIAL set to 3
      • ACM20 = 127.3 Watts
      • PID Reaches = 233°C
    4. At 9:30 PM
      • DIAL set to 4.
      • ACM20 = 206 Watts
      • PID Reaches = 387°C
    5. At 10:30 PM
      • DIAL set to 5
      • ACM20 = 269 Watts
      • PID Reaches = 506°C
    6. At 11:30 PM (ish?!)
      • DIAL set to 6
      • ACM20 = 337 Watts
      • PID Reaches = 652°C
    7. At 12:55 (seems I fell asleep) P.T.O.
      • DIAL set to 7
      • ACM20 = 424.5 Watts
      • PID Reaches = 788°C
    8. At 2 AM (snoring again)
      • DIAL set to Max (7.7 opex)
      • ACM20 = 462.8 Watts
      • PID Reaches = 831°C
      • Ambient 27.9 in EX Room

    Before shutting Down I place a S-Type Platinium Rhodium Thermocouple in the LHS Port it reads a temp of 1050°C - 831 = 219°C


    Remember in this test as in the Main Run there are Short Thermocouples in a case so barely the Tip is in the Thermowell. But as yet N° ceramic wool has been added. The Tips are Naked and yet there is a 219°C difference LESS than the True Temperature.


    When the Wad of ceramic Wool is added in the Main Test Run to protect the thermocouples from (ss) so I can at least gather Data the (False) Temperature from the Real or True Temperature will be even greater. This is of course very significant and a must be bourne in mind when estimating energy in versus energy out Though thankfully there are other clear indications.


    Energy-in of course is known.

    Main Test Run

    Wednesday November 7th 2018

    • LHP + RHP Thermos are short 5cm
    • ceramic Wool (Bob will find on take Down)
    • LION Tubes
    • Inside
      • LHP = Tube = D2 Pads + Silver (leaf)
      • RHP = Tube = D2 Pads + Stainless Steel
    • Outside
      • LHP = Tube x7 Windings of Cu Foil + ogivei
      • LHP = Tube x7 Windings of Cu Foil + Ifgubie
    • PID Thermocouple is connected to Back RH Post Stainless Steel
    • TI = Thermocouple is from Front New Hole Top o
    • Tg = is connected to LH Post at rear-side-Tube

    Test Started at 11.15 pm.

    1. At 11.15 P.M (Wed 7.1.2018)
      • DIAL set to 1
      • ACM20 = 14.2 Watts
      • PID Reaches = 32°C
      • Ambient is 25 in EX Room
      • Also a 1.4 setting
    2. 5.00 Am Thursday Morning
      • DIAL set to 2
      • ACM20 =
      • PID Reaches = 37°C
      • Also a 2.1 set
      • 2.3 set
      • 2.5 set
      • 2.7 set
    3. 12:30 PM Thursday Night.
      • DIAL Set to 3
      • Also a 3.5 set
      • ACM20 = 161 Watts
      • PID Reaches = 295°C
    4. Friday 5 PM
      • DIAL set to 4
      • ACM20 = 170 Watts
      • PID Reaches = 410°C
      • Also a 4.5 set
    5. Saturday 11.15 am
      • DIAL set to 5
      • ACM20 = 193 Watts
      • PID Reaches = 450°C
      • Also a 5.5 set
    6. Saturday 7.28 PM
      • DIAL set to 6
      • ACM20 = 343 Watts
      • PID Reaches = 594°C
      • Also a 6.5 set
    7. Sunday 8PM
      • DIAL set to 7
      • ACM20 = 420 Watts
      • PID Reaches 700°C
      • Ambient 27.5 + increasing in EX Room. See SD card for Details

    Start of Thermal Shock Process

    DIAL Turned back down to (6)

    Temp falls to 491-592°C

    ACM20 = 334.8 Watts.

    Monday 1.57 pm DIAL set to 7.7
    Tuesday 1.58 am DIAL turned back down to (6)
    " 3.am PID stable at 597°C
    " 12.20am PID still at 597°C
    • N° Pulsing-steady state
    • Watts = 332.7
    • Tuesday
      • 1.39pm PID up 597°C-598°C
    • Wednesday
      • 2.53am PID steady at 598°C
    • Watts = 331.9
      • up +6°C from first Thermal Shock

    Wednesday 7.30pm

    Put DIAL back up to 7.7 from 6

    This turned out to be a serious error

    Clearly it should have been raised slowly until the New Fire is stable it is very easy to Put it out. Patience is Required to Kindle this FIRE.

    At about 8.30 something broke.

    1. Screeching sound
    2. Arks and Sparks
    3. Sparkler sound for approx 30 seconds

    At 10.45 PM PID Temp still at 60°C so over 24 hours for cool down.

    • Slow, careful Multi Tube approach

    NEXT.-AIM = PULL the Plug Out. Keep it Simple.

    With PATIENCE I believe that this is a doable project.


    Empty Tube + Copper x7 @ Wire
    Calibration Run DIAL Test Run Difference.
    13.7 Watts 1 14.2 Watts +.5 Watts
    PID 33°C 32°C -1 °C
    18.28 Watts 2 ?
    PID 111°C
    127.3 Watts 3 161 Watts +33.7 Watts
    PID 233°C 295°C +62 °C
    206 Watts 4 170 Watts - 36 Watts
    PID 387 °C 410°C + 23 °C
    269 Watts 5 193 Watts -76 Watts
    PID 506°C 450 °C - 56 °C
    337 Watts 6 343 Watts +6 Watts
    PID 652 °C 619 °C -33 °C
    424.5 Watts 7 420 Watts -4.5 Watts
    PID 778 °C 700 °C -78 °C


    Simple liste for Like comparison NOT allowing for a correcting False Temp due to Set up.

  • I asked LION what PID settings he employed in his recent experiments. He has answered here http://disq.us/p/1xvsphz.


    I take it that the 1,300 C setting he refers to is the set point temperature. Since both thermocouples were below 900 C for the entire duration of these experiments I expect that the PID was always outputting its maximal control signal, whatever that was. This would be true for the control as well as the active runs.


  • ...I suspect that there is an extrinsic signal that is being impressed on both the current and thermocouple channels from somewhere.


    At +20°C (relative to the cold junction) a Type K thermocouple's voltage will be around 800 µV. The rate constant is thus around 40µV/°C. So the jitter in the temperature data represents about 80 microvolts p-p at the DAQ input


    I suspect that this noise in the therm-A trace may be caused by leakage from the DC chopper (PWM) power controller. LION's post seems to confirm that he had the PWM connected in his setup. See my previous post for a sample waveform and harmonic spectrum.

  • I suspect that this noise in the therm-A trace may be caused by leakage from the DC chopper (PWM) power controller. LION's post seems to confirm that he had the PWM connected in his setup. See my previous post for a sample waveform and harmonic spectrum.


    This seems right to me. The data acquisition rate is 1 Hz here which is far lower than the frequencies in the PWM output. I think that the periodic signal (with period ~23 sec) seen in the ThermoA trace is a severely aliased version of the high-frequency PWM output.

  • I don't think I am looking at an artefact of plotting. You and I seem to be dealing with datasets that are actually different. For instance, in the 53 seconds between 7:52:73 and 7:53:36 my datasheet only shows 24 valid ThermoA data values (the rest show as "Error"). In your plot, on the other hand, there are 53 separate values. This leads me to think that your plot is showing data where interpolated values have been imputed for about half of the data points.

    My data came from this post:

    LION-AG Experiment


    I downloaded the zip file called

    LION4_noholes_interpolated.zip


    and then plotted the data from 003.csv


    Where did your data come from?

  • Robert Horst


    Bob Greenyer obtained the files from LION and has made them available for download at https://goo.gl/8hvrrM. This is where I got the data I am using. Can must have obtained his data from the same place.


    The ThermoA and ThermoB columns in the file (LOG00002.csv) for the active run contain a large number of cells with error messages. Can mentions that when he downloaded the file he filled these cells with interpolated values. I have not interpolated any values. I have just left the cells blank.


    P.S. The data files I have been working with are available (for the next 3 weeks) at https://ufile.io/fwewc. In generating these files I took the original LION data files posted by Bob Greenyer, stripped out blank lines, then replaced every "Error" entry by a blank. The "error" entries are profuse on the thermoA and B channels and make up about 1/4 of all those entries.

  • I have plotted the temperature and power data found in the table in LION's notes (thanks to can for transcribing the notes).




    The blue markers show data from a calibration run in which the heating chamber of the Looking For Heat device was left empty. The orange markers show data from the first 4 days of LION's experiment using a fueled reactor in the heating chamber. The green square come from a later era of the same experiment after LION put the system through what he called a "thermal shock".


    I am not completely sure where the temperature data come from. In particular, i am not sure how they would correspond to the temperatures acquired and stored on the data card that LION supplied. Are the 2 orange data points at input powers just below 200 Watts supposed to be illustrating excess heat?

  • Bruce__H

    I'm glad you appreciated the effort, but please also check the original document to make sure that the values are correct, because I transcribed it in 30 minutes or so and it likely contains errors and certainly a few placeholder words where I couldn't figure out LION's writing at the time of reading.